# d2Cache: Accelerating Diffusion-Based LLMs via Dual Adaptive Caching

### v1

@@ C.2 More visualization results on attention rollout
C.3 More visualization results on KV state dynamics
To further substantiate our findings in Section 3.2, we visualize additional KV state dynamics. In Figure 8, which visualizes the trajectories of the key states and value states of the same masked token during decoding, both are closely aligned in both trajectory shape and magnitude, and both exhibit the same gradual–rapid–stable dynamic pattern. This result suggests that, for both key states and value states, it is sufficient to update them only during the rapid-change phase, where these KV states can be safely cached for reuse during the other two phases.
We hypothesize that this rapid change arises because tokens are particularly sensitive to changes in their local context. Specifically, at step $t$, if a masked token $x_{t}^{i}$ is located near another masked token $x_{t}^{j}$ that is decoded, then at step $t+1$ the embedding of $x_{t}^{j}$ changes from [MASK] to the embedding of a concrete token. This provides $x_{t}^{i}$ with additional contextual information; the smaller the distance $\lvert i-j\rvert$, the more tightly constrained the context becomes, thereby substantially altering the model’s representation of $x_{t}^{i}$. These observations motivate the introduction of distance-aware decay into the certainty density, as defined in Equation 3.
@@ C.3 More visualization results on KV state dynamics
Figure 9: Ablation studies on hyperparameters (a) $p$ and $k$, and (b) $\sigma$. All experiments are conducted on Dream-Inst with the HumanEval dataset.
C.4 Ablation on hyperparameters
To determine the optimal hyperparameters, we conduct systematic experiments on Dream-Inst using the HumanEval dataset. As shown in Figure 9 (a), the number of masked tokens selected per step has a substantial influence on performance. While larger values of $k$ generally improve performance, increasing $k$ beyond 32 yields diminishing returns and may even result in slight degradation. This suggests that $k=32$ provides the most robust performance across different values of $p$. In parallel, the cumulative probability threshold $p$ controls the proportion of probability mass retained among the remaining tokens at each step and primarily affects throughput. The performance for $p=0.10$ and $p=0.20$ is nearly indistinguishable, indicating that the marginal benefit of increasing $p$ becomes negligible beyond 0.1. Considering the trade-off between efficiency and accuracy, we therefore set $p=0.1$ and $k=32$ throughout our experiments unless explicitly specified.
The standard deviation $\sigma$ of the Gaussian function in Equation 3 determines the range considered during certainty prior selection. A smaller $\sigma$ biases the selection of masked tokens for updating and decoding towards those closer to the prompt or already decoded tokens, whereas a larger $\sigma$ incorporates a broader local context. While the influence of $\sigma$ on decoding order has been discussed in Section 5.3, we further investigate its impact on performance using Dream-Inst. As shown in Figure 9 (b), when $\sigma=1$, Dream-Inst tends to decode in an approximately autoregressive manner, leading to degraded performance. Conversely, when $\sigma=80$, resembling non-autoregressive decoding, Dream-Inst prematurely generates the [EOS] token, which also degrades performance. These observations are consistent with the properties of LLaDA (Nie et al., 2025). We find that $\sigma=10$ yields the best performance, enabling a flexible, quasi left-to-right decoding order.

### v2

@@ E.1 Experimental results on the Base variants
E.2 Experimental results under parallel decoding settings
To enable a fair comparison across all methods and to verify the generalizability of d2Cache under alternative decoding strategies, we evaluate all approaches using the parallel decoding strategy, where the threshold is set to 0.9 following Wu et al. (2025). As shown in Table 6, our method achieves up to 48.3× acceleration over the single-token-per-step baseline while maintaining performance comparable to all other baselines, which clearly demonstrates the broad applicability of d2Cache.
E.3 Experimental results under long-context settings
To further assess our method’s performance under long-context settings, we further evaluate our method on Dream-Inst under with a longer generation length 1024. As shown in Table 7, we observe that other methods—due to their coarse-grained nature—experience severely degraded acceleration as the context length increases. In contrast, d2Cache maintains substantial speedups without performance loss even in long-context scenarios, owing to its fine-grained two-stage token selection. These results demonstrate that d2Cache also performs well in long contexts.
E.4 More visualization results on attention rollout
E.5 More visualization results on KV state dynamics
To substantiate our findings in Section 3.2, we visualize additional KV state dynamics. In Figure 9, which visualizes the trajectories of the key and value states of the same masked token during decoding, both are closely aligned in both trajectory shape and magnitude, and both exhibit the same gradual–rapid–stable dynamic pattern. This result suggests that, for both key and value states, it is sufficient to update them only during the rapid-change phase, where these KV states can be safely cached for reuse during the other two phases. We hypothesize that this rapid change arises because tokens are particularly sensitive to changes in their local context. Specifically, at step $t$, if a masked token $x_{t}^{i}$ is located near another masked token $x_{t}^{j}$ that is decoded, then at step $t+1$ the embedding of $x_{t}^{j}$ changes from [MASK] to the embedding of a concrete token. This provides $x_{t}^{i}$ with additional contextual information; the smaller the distance $\lvert i-j\rvert$, the more tightly constrained the context becomes, thereby substantially altering the model’s representation of $x_{t}^{i}$. These observations motivate the introduction of distance-aware decay into the certainty density, as defined in Equation 3.